carrier In addition, Gold E cost if the product is lost in transit service to be $90 per unit. Brinks picked up the goid bars from Gold Examiner on March 30, and delivery to the bank occurred on 1. How many Ree Reg 2 to4
Req 2 to 4 in the first field.) Cnet Ciear entry
in the first account field.) Journal entry worksheet any necessary entry when Brinks has picked up the gold bars from Gold Examiner 30, 2018 Clear entry
n list Journal entry worksheet 2 Record any necessary entry upon delivery of the gold bars to the bank. ral Jou DebitCredit April 01, 2018 Record entry Clear entry Prey11 of 20
Answer #1
1
Number of performance obligations = 2
Delivery of gold and additional insurance
2
March 01, 2018 Cash 160000
      Deferred revenue—gold bars 150400
      Deferred revenue—insurance 9600
March 30, 2018 Deferred revenue—gold bars 150400
       Sales revenue 150400
April 01, 2018 Deferred revenue—insurance 9600
     Service revenue 9600
Workings:
Value of the gold bars 132540 =94*1410
Stand­alone selling price of the insurance 8460 =94*90
Total of stand­alone prices 141000
Deferred revenue—gold bars 150400 =160000*132540/141000
Deferred revenue—insurance 9600 =160000*8460/141000